# FLAME: A Federated Learning Benchmark for Robotic Manipulation

### v1

@@ Abstract
Recent progress in robotic manipulation has been fueled by large-scale datasets collected across diverse environments. Training robotic manipulation policies on these datasets is traditionally performed in a centralized manner, raising concerns regarding scalability, adaptability, and data privacy. While federated learning enables decentralized, privacy-preserving training, its application to robotic manipulation remains largely unexplored. We introduce FLAME (Federated Learning Across Manipulation Environments), the first benchmark designed for federated learning in robotic manipulation. FLAME consists of: (i) a set of large-scale datasets of over 160,000 expert demonstrations of multiple manipulation tasks, collected across a wide range of simulated environments; (ii) a training and evaluation framework for robotic policy learning in a federated setting. We evaluate standard federated learning algorithms in FLAME, showing their potential for distributed policy learning and highlighting key challenges. Our benchmark establishes a foundation for scalable, adaptive, and privacy-aware robotic learning.
@@ I INTRODUCTION
[…]
@@
In summary, our contributions are as follows:
  • FLAME, a federated learning benchmark for robotic manipulation, consisting of a large-scale dataset collected across multiple simulated environments and tasks, and a federated learning and evaluation framework designed for robotic manipulation.
  • A rigorous evaluation of different federated learning baselines, demonstrating the need for more research for federated approaches for robotic manipulation training in a distributed setting.
  • To the best of our knowledge, this work presents the first application of federated learning to robotic manipulation strategies.
We believe our benchmark will serve as a valuable resource for the robotics and machine learning communities, fostering advancements in robotic manipulation. The code and datasets of FLAME will be made publicly available upon acceptance of the paper.
@@ IV EVALUATION
(a) Offline Evaluation: RMSE ×$\times 10^{-2}(\downarrow)$
@@
FedAvg $0.24\pm 0.43$ $0.84\pm 0.37$ $0\pm 0$ $0\pm 0$
FedAvgM $0\pm 0$ $0.54\pm 0.49$ $0\pm 0$ $0\pm 0$
FedOpt $0.28\pm 0.45$ $0.70\pm 0.46$ $0\pm 0$ $0\pm 0$
Krum $0.10\pm 0.30$ $0.86\pm 0.38$ $0\pm 0$ $0\pm 0$
(b) Online Evaluation: Normalized Success Rate ()($\uparrow$)( ↑ )
@@ IV-A Experimental Setup
[…]
Local Imitation Learning Agent: For each environment, we instantiate the local agent as a multi-modal neural network that integrates visual and low-dimensional state inputs to predict continuous control actions. The architecture employs a CNN encoder for processing RGB images of size (64, 64, 3), consisting of three convolutional layers with 32, 64, and 128 filters, respectively, each with a 3×3 kernel, stride of 2, and ReLU activation, followed by a fully connected layer of 256 units. In parallel, an MLP encoder processes the low-dimensional state input through two fully connected layers with 256 and 128 units, both using ReLU activations. The extracted visual and state features are concatenated into a 384-dimensional latent representation and passed through a policy network comprising two fully connected layers of 512 and 4 units, where the final layer applies a Tanh activation to constrain the action outputs to the range [-1, 1]. The model is trained using supervised learning with a Mean Squared Error (MSE) loss, optimizing the action predictions against expert demonstrations using the Adam optimizer with a learning rate of 1e-4. To enable decentralized training, the agent is deployed within a federated learning framework. To incorporate updates from multiple distributed environments, we benchmark different aggregation strategies, as described in the following section.
@@
Training Setup: We trained all federated learning models for a total of $30$ aggregation rounds, with each client performing $50$ local training epochs per round. The number of available clients was set to $400$, corresponding to the number of distinct training environments. Due to computational constraints, $20$ random clients were selected per training round. Model validation and testing were conducted on $20$ previously unseen environments, with the best-performing model on the evaluation set subsequently assessed on the test set. All models were trained on a system equipped with a single NVIDIA A100100100100 GPU, $64$ CPU cores, and $250$ GB of allocated memory.

### v2

@@ Abstract
Recent progress in robotic manipulation has been fueled by large-scale datasets collected across diverse environments. Training robotic manipulation policies on these datasets is traditionally performed in a centralized manner, raising concerns regarding scalability, adaptability, and data privacy. While federated learning enables decentralized, privacy-preserving training, its application to robotic manipulation remains largely unexplored. We introduce FLAME (Federated Learning Across Manipulation Environments), the first benchmark designed for federated learning in robotic manipulation. FLAME consists of: (i) a set of large-scale datasets of over 160,000 expert demonstrations of multiple manipulation tasks, collected across a wide range of simulated environments; (ii) a training and evaluation framework for robotic policy learning in a federated setting. We evaluate standard federated learning algorithms in FLAME, showing their potential for distributed policy learning and highlighting key challenges. Our benchmark establishes a foundation for scalable, adaptive, and privacy-aware robotic learning. The code is publicly available at https://github.com/KTH-RPL/ELSA-Robotics-Challenge.
@@ I INTRODUCTION
To address this gap, we introduce FLAME (Federated Learning Across Manipulation Environments), a benchmark designed to evaluate federated learning strategies for robotic manipulation tasks. Our benchmark provides a set of large-scale datasets of diverse manipulation tasks collected across multiple settings, incorporating variations in lighting conditions, textures, object appearances, and camera viewpoints, encompassing over $15\text{M}$ data samples. Furthermore, FLAME integrates these distributed datasets into a FL framework, enabling a rigorous evaluation of FL algorithms and serving as a foundation for developing more robust federated learning manipulation strategies.
@@
In summary, our contributions are as follows: (i) FLAME, a federated learning benchmark for robotic manipulation, consisting of a large-scale dataset collected across multiple simulated environments and tasks, and a federated learning and evaluation framework designed for robotic manipulation; (ii) a rigorous evaluation of different federated learning baselines, demonstrating the need for more research for federated approaches for robotic manipulation training in a distributed setting; (iii) to the best of our knowledge, this work presents the first application of federated learning to robotic manipulation strategies
We believe our benchmark will serve as a valuable resource for the robotics and machine learning communities, fostering advancements in robotic manipulation. The code and datasets of FLAME are publicly available at https://github.com/KTH-RPL/ELSA-Robotics-Challenge.
@@ IV EVALUATION
(a) Offline Evaluation: RMSE $\times 10^{-2}(\downarrow)$
@@
FedAvg $0.24\pm 0.43$ $84\pm 0.37 0±0$$\pm 0$ $0\pm 0$
FedAvgM $0\pm 0$ $54\pm 0.49 0±0$$\pm 0$ $0\pm 0$
FedOpt $0.28\pm 0.45$ $70\pm 0.46 0±0$$\pm 0$ $0\pm 0$
Krum $0.10\pm 0.30$ $86\pm 0.38 0±0$$\pm 0$ $0\pm 0$
(b) Online Evaluation: Normalized Success Rate $(\uparrow)$
@@ IV-A Experimental Setup
Tasks: For our evaluation, we select four distinct tasks: $Q=\{$Slide block to target (Slide Block), Close Box (Close box), Insert onto Square Peg (Peg in Square), Scoop with Spatula (Scoop)}\}. In each of these environments, the robot’s action space is composed of joint velocities as well as a binary command to open or close the gripper. Meanwhile, the observation space includes the robot’s joint positions and a single-view camera feed that supplies image observations. This setup allows us to capture both low-dimensional kinematic information and high-dimensional visual context. For each of these tasks, we follow the data collection procedure described in the previous section, containing $N=420$ different environments and $K=100$ demonstrations per environment, of which $400$ environments are used for training, $10$ for validation, and $10$ for testing. We highlight the selected factors of variation in Table I.
Local Imitation Learning Agent: For each environment, we instantiate the local agent as a multi-modal neural network that integrates visual and low-dimensional state inputs to predict continuous control actions. The architecture employs a CNN encoder for processing RGB images of size (128, 128, 3), consisting of three convolutional layers with 32, 64, and 128 filters, respectively, each with a 3×3 kernel, stride of 2, and ReLU activation, followed by a fully connected layer of 256 units. In parallel, an MLP encoder processes the low-dimensional state input through two fully connected layers with 256 and 128 units, both using ReLU activations. The extracted visual and state features are concatenated into a 384-dimensional latent representation and passed through a policy network comprising two fully connected layers of 512 and 4 units, where the final layer applies a Tanh activation to constrain the action outputs to the range [-1, 1]. The model is trained using supervised learning with a Mean Squared Error (MSE) loss, optimizing the action predictions against expert demonstrations using the Adam optimizer with a learning rate of 1e-4. To enable decentralized training, the agent is deployed within a federated learning framework. To incorporate updates from multiple distributed environments, we benchmark different aggregation strategies, as described in the following section.
@@
Training Setup: We trained all federated learning models for a total of $30$ aggregation rounds, with each client performing $50$ local training epochs per round. The number of available clients was set to $400$, corresponding to the number of distinct training environments. Due to computational constraints, $20$ random clients were selected per training round. Model validation and testing were conducted on $20$ previously unseen environments, with the best-performing model on the evaluation set subsequently assessed on the test set. All models were trained on a system equipped with a single NVIDIA A100100 GPU, $64$ CPU cores, and $250$ GB of allocated memory.